The following data were taken from the records of Clarkson Company for the fiscal year ended June 30,2014.

Raw Materials

Inventory 7/1/13 $48,000

Factory Insurance $4,600

Raw Materials

Inventory 6/30/2014 $39,600

Factory Machinery

Depreciation $16,000

Finished Goods

Inventory 7/1/13 $96,000

Factory Utilities $27,600

Finished Goods

Inventory 6/30/14 $75,900

Office Utilities Expense $8,650

Work in Progress

Inventory 7/1/13 $19,800

Sales Revenue $534,000

Work In Progress

Inventory 6/30/14 $18,600

Sales Discounts $4,200
Direct Labor $139,250 Plant Managers Salary $58,000
Indirect Labor $24,460 Factory Property Taxes $9,600
Accounts Receivable $27,000 Factory Repairs $1,400
Raw Materials Purchases $96,400
Cash $32,000

A) Prepare a cost of goods manufactured schedule (Assume all raw materials used were direct materials).

B) Prepare an income statement through gross profit

C)Prepare the current assets section of the balance sheet at June 30, 2014
Answer #1
Concepts and reason

Income statement: All companies, to calculate the net profit or net loss prepare income statement. The statement of income is one of the important financial statements. It is prepared by deducting all revenue expenses from total revenue receipt during that period for which statement is prepared.

Statement of Owner’s Equity: It is a financial statement showing the changes in the equity section during the financial year. Owner’s Equity is also known as Shareholder’s Equity. It represents the capital contributed, withdrawals, dividends and the net income earned during the year.

Balance sheet: It is the financial statement which includes the assets, equity, and liabilities of a business at a particular time. It shows the company’s financial position. The steps for preparations of Balance Sheet are as follows:

1. The first heading in the preparation of balance sheet is assets. Under which the balances of currents assets are presented first, then the balances of long term investments are entered and then the balances of plant, property, and equipment are entered.

2. The second heading in the balance sheet preparation is shareholder’s equity and liabilities. The liabilities are presented first, under which the balances of current liabilities and then the long- term debts balances are presented. Stockholder’s equity balances are entered after the liabilities.

3. Ensure that the balances of both the assets portion and liabilities and shareholder’s equity portion are equal.

Fundamentals

Inventory: Inventory refers to the material or goods that are seized by the organizations with a motive of selling it after some processing. It includes various categories that are finished goods, work in process, and raw material.

Work-in-progress inventory (WIP): It is the raw materials, overhead costs, and labor costs incurred for products at various stages of the process of production. It is the part of the inventory asset account of the balance sheet.

Accounts Receivable: It is an asset for the company. It is a consideration, which the company has the right to receive the goods or services provided to the customers. It is the payment which the company will receive in the future from its customers.

The cost of goods manufactured (COGM): It is the schedule that shows the calculation of the cost of production that was produced in the accounting period. It consists of all the costs that have been incurred for the production.

Cost of goods sold (COGS): These are the direct costs that are attributed to the production of the goods that have been sold by a company. It includes the labor cost incurred to produce the goods and the cost of the materials used. It does not include the indirect expenses.

Current assets: These are the accounts of the balance sheet that are likely to be converted into cash in a year. It includes cash, inventory, accounts receivables, securities that are marketable, prepaid expenses, and other assets that are likely to be converted into cash in a year.

Stockholder’s equity: Shareholder’s Equity is also known as owner’s equity. It is a part of Balance sheet other than assets and liabilities. It represents the stake on equity by the investors in the firm.

(A)

Prepare a schedule for COGM:

A
C Company
Cost of goods manufactured schedule
for the year end June 30, 2014
2
Amount (S) Amount (S)
3 Work in process July19 Total manufacturing overheads
20 Total manufacturing costs
21 Total cost of work in process
22 Less: Work in process June

Working notes:

Compute the total raw material available for use:

RawmaterialAvailableforuse=RawmaterialinventoryJuly1,2013+Rawmaterialpurchases=$48,000+$96,400=$144,400\begin{array}{c}\\{\rm{Raw materia}}{{\rm{l}}_{{\rm{Available for use}}}} = {\rm{Raw material inventor}}{{\rm{y}}_{{\rm{July 1, 2013}}}} + {\rm{Raw material purchases}}\\\\ = \$ 48,000 + \$ 96,400\\\\ = \$ 144,400\\\end{array}

Compute the amount of direct materials used:

Directmaterialsused=TotalrawmaterialAvailableforuseRawmaterialinventoryJune31,2014=$144,400$39,600=$104,800\begin{array}{c}\\{\rm{Direct materials used}} = {\rm{Total raw materia}}{{\rm{l}}_{{\rm{Available for use}}}} - {\rm{Raw material inventor}}{{\rm{y}}_{{\rm{June 31, 2014}}}}\\\\ = \$ 144,400 - \$ 39,600\\\\ = \$ 104,800\\\end{array}

Compute the amount of total manufacturing overheads:

Manufacturingoverheads=Plantmanagerssalary+Indirectlabor+Factoryutilities+Factorymachinedepreciation+Factoryinsurance+Factorypropertytaxes+Factoryrepairs=$58,000+$24,460+$27,600+$16,000+$4,600+$9,600+$1,400=$141,660\begin{array}{c}\\{\rm{Manufacturing overheads}} = {\rm{Plant managers' salary}} + {\rm{Indirect labor}}\\\\ + {\rm{Factory utilities}} + {\rm{Factory machine depreciation}}\\\\ + {\rm{Factory insurance}} + {\rm{Factory property taxes}}\\\\ + {\rm{Factory repairs}}\\\\ = \$ 58,000 + \$ 24,460 + \$ 27,600\\\\ + \$ 16,000 + \$ 4,600 + \$ 9,600 + \$ 1,400\\\\ = \$ 141,660\\\end{array}

Compute the total manufacturing costs:

Totalmanufacturingcosts=Manufacturingoverheads+Directlabor+Directmaterialsused=$104,800+$139,250+$141,660=$385,710\begin{array}{c}\\{\rm{Total manufacturing costs}} = {\rm{Manufacturing overheads}} + {\rm{Direct labor}}\\\\ + {\rm{Direct materials used}}\\\\ = \$ 104,800 + \$ 139,250 + \$ 141,660\\\\ = \$ 385,710\\\end{array}

Compute the total cost of work in process:

Totalcostsofworkinprocess=WorkinprocessJuly1,2013+Totalmanufacturingcosts=$19,800+$385,710=$405,510\begin{array}{c}\\{\rm{Total costs of work in process}} = {\rm{Work in proces}}{{\rm{s}}_{{\rm{July 1, 2013}}}} + {\rm{ Total manufacturing costs}}\\\\ = \$ 19,800 + \$ 385,710\\\\ = \$ 405,510\\\end{array}

Compute the cost of goods manufactured:

Costofgoodsmanufactured=TotalcostofworkinprocessWorkinprocessjune30,2014=$405,510$18,600=$386,910\begin{array}{c}\\{\rm{Cost of goods manufactured}} = {\rm{Total cost of work in process}} - {\rm{Work in proces}}{{\rm{s}}_{{\rm{june 30, 2014}}}}\\\\ = \$ 405,510 - \$ 18,600\\\\ = \$ 386,910\\\end{array}

(B)

Prepare the statement of income for the year ended June 30, 2014:

25
26
(4,200)
A
C Company
Income statement (Partial)
for the year end June 30, 2014
Amount () Amount (S)
27 Sales revenue:
28

Working notes:

Compute the amount of net sales:

Netsales=SalesrevenueSalesdiscount=$534,000$4,200=$529,800\begin{array}{c}\\{\rm{Net sales}} = {\rm{Sales revenue}} - {\rm{Sales discount}}\\\\ = \$ 534,000 - \$ 4,200\\\\ = \$ 529,800\\\end{array}

Compute the amount of cost of goods available for sale:

Costofgoodsavailableforsale=FinishedgoodsinventoryBeginning+Costofgoodsmanufactured=$96,000+$386,910=$482,910\begin{array}{c}\\{\rm{Cost of goods available for sale}} = {\rm{Finished goods inventor}}{{\rm{y}}_{{\rm{Beginning}}}}\\\\ + {\rm{Cost of goods manufactured}}\\\\ = \$ 96,000 + \$ 386,910\\\\ = \$ 482,910\\\end{array}

Compute the amount of cost of goods sold:

Costofgoodssold=CostofgoodsavailableforsaleFinishedgoodsAttheend=$482,910$75,900=$407,010\begin{array}{c}\\{\rm{Cost of goods sold}} = {\rm{Cost of goods available for sale}} - {\rm{Finished good}}{{\rm{s}}_{{\rm{At the end}}}}\\\\ = \$ 482,910 - \$ 75,900\\\\ = \$ 407,010\\\end{array}

Compute the amount of gross profits:

Grossprofit=NetsalesCostofgoodssold=$529,800$407,010=$122,790\begin{array}{c}\\{\rm{Gross profit}} = {\rm{Net sales}} - {\rm{Cost of goods sold}}\\\\ = \$ 529,800 - \$ 407,010\\\\ = \$ 122,790\\\end{array}

(C)

Prepare the current assets portion of the balance sheet:

C Company
Balance sheet (Partial)
June 30, 2014
Amount (S) Amount (S)
39
40
32.000
27.000
41 Current Assets
42 Cash
43 Accoun

Working notes:

Compute the amount of inventories:

Inventory=Finishedgoods+Workinprogress+Rawmaterial=$75,900+$18,600+$39,600=$134,100\begin{array}{c}\\{\rm{Inventory}} = {\rm{Finished goods}} + {\rm{Work in progress}} + {\rm{Raw material}}\\\\ = \$ 75,900 + \$ 18,600 + \$ 39,600\\\\ = \$ 134,100\\\end{array}

Compute the amount of total current assets:

Totalcurrentassets=Cash+Accountsreceivale+Inventories=$32,000+$27,000+$134,100=$193,100\begin{array}{c}\\{\rm{Total current assets}} = {\rm{Cash}} + {\rm{Accounts receivale}} + {\rm{Inventories}}\\\\ = \$ 32,000 + \$ 27,000 + \$ 134,100\\\\ = \$ 193,100\\\end{array}

Ans: Part A

The cost of goods manufactured for the year is $386,910.

Part B

The gross profit for the year is $122,790.

Part C

The total of current assets in the balance sheet is $193,100.